# Towards Reducible Uncertainty Modeling for Reliable Large Language Model Agents

### v1

@@ 3.2 A Unified View on Existing UQ Setups
We further draw a connection between this trajectory-level UQ problem and the process reward modeling problem [uesato2022solving, lightman2024lets, li2025pqm] where we assign credit (reward) for each intermediate reasoning step, and then it can be utilized to evaluate the reasoning trajectory by aggregating the rewards across steps to derive a trajectory-level reward. For example, lightman2024lets proposed the product of the step-wise probabilities and the minimum probability between steps as the aggregation method. These are analogs to Eq. 3 and Eq. 4, respectively.
Figure 3: Limitation of existing UQ (a) and our suggestion (b). Prior works just concern the evidentiality when designing or evaluating UQ while neglecting interactivity, so they may fail to reliably capture the agent’s failure. We urge expanding a dimension of interest, interactivity, and moving on to reducible uncertainty modeling of agents.
Limitation of existing frameworks.
From this unified view, we see that they all model the multi-step uncertainty as a uni-directional propagation, i.e., every step-level uncertainty positively contributes to the total uncertainty, without distinguishing action types at each turn. One might claim that the average uncertainty $\frac{1}{T}\sum^{T}_{t=1}U(\mathcal{F}_{t}|\mathcal{F}_{t-1})$ can already be fine, as it might produce a lower value for the chain of certain actions and vice versa. However, it does not consider the type of action which is critical in agentic setups, where an agent’s interactivity is key to earning a reward. See Fig. 3 (a), a scenario wherein a user does not fully specify their goal upfront. After searching for options, the agent either reasons itself or interacts with the user to delegate decisions. Although both trajectories are evidential, interaction-oriented ones tend to achieve higher rewards, despite reasoning-oriented ones often yielding greater confidence. Yet, existing methods can produce misleading uncertainty estimates, fail to meet the desiderata (Eq 1).
4 How to Model Uncertainty Dynamics of Interactive Agents?
4.1 Unique Challenge
In contrast to the LLM UQ, agent UQ spans a multi-turn trajectory in an open environment. This presents a unique challenge: modeling cascade uncertainty in an interactive system. Here, the desired UQ method should be able to precisely aggregate uncertainties across the entire trajectory and account for additional information gained through interactions. While there is emerging research on step-wise uncertainty aggregation [kang2025scalable, fu2025deep], they examine an isolated, noninteractive setting where an agent does not receive any external feedback during inference (e.g., Fig. 1, b). Meanwhile, agent UQ considers a setup full of interactions with a user and an environment, making it a non-isolated, interactive system.
Since an agent in this open world can gather additional information, make clarifications, and request confirmations at each step by interacting with the user and external databases through conversations and tool calls, the uncertainty dynamics in this interactive system includes not only propagations of turn-level uncertainty, but also reductions. As shown in Fig. 3, the existing perspective does not consider “interactivity” in agent UQ and prohibits reducible uncertainty modeling. In this paper, we argue that we must introduce interactivity in agent UQ to model bi-directional uncertainty flow.
4.2 Agent UQ as a Conditional Uncertainty Reduction Process
We need a new perspective to model an agent’s uncertainty dynamics that accounts for both reductions and increases, depending on the agent’s interactivity at each step. While there are multiple ways to realize this concept, we propose one possible implementation as a prototype: information gating, which goes hand in hand with the information-theoretic measure of uncertainty noted in Sec. 3.
Specifically, we derive a lower bound of total uncertainty $U(\mathcal{F}_{\leq T})$ that admits a selective reduction or increase in uncertainty throughout the trajectory:
Action Category Interactivity Example
Table 1: Example action classes in Airline booking scenario in $\tau$-bench. A number of the agent’s actions can be categorized into five classes, which may reduce or increase the total uncertainty. See Appendix E for extended analyses.
Lemma 1 (Extrema of Information Gating).
Proof.
If $A_{i}\in\mathcal{A}^{-}$ (resp. $A_{i}\notin\mathcal{A}^{-}$) for all $1\leq i\leq T-1$, then $\tilde{U}(\mathcal{F}_{\leq T})$ becomes a monotonic uncertainty reduction (resp. propagation) process, deriving above inequalities (See Appendix F for details). ∎
Interpretation.
Lemma 1 states that an agent’s total uncertainty can be expressed as an ever-decreasing process starting from a task-specific initial uncertainty reduced by the amount of extra information delivered by current observation, given the previous trajectory till its final behavior in the ideal case; whereas it is continually stacked up by the amount of joint entropy between action and observation per turn in worst case. Together with the directional interpretation turn-level uncertainty (increase or decrease), these analytic extrema help us better understand the numerical uncertainty, which is usually hard to interpret [van2019communicating, belem2024perceptions, kirchhof2025self].
Sketch of implementation.
To operationalize this idea, we first need an action classifier to determine whether $A_{i}\in\mathcal{A}^{-}$ or not. One can adopt a model-based evaluator [kim2025beyond] and a rule-based verifier [zeng2025reinforcing] to assess the interactivity and evidentiality of actions, respectively, where $A_{i}\in\mathcal{A}^{-}$ if and only if $A_{i}$ is both evidential and interactive. An example categorization of actions in an airline assistant task is provided in Table 1. After that, four different quantities need to be estimated: (1) initial query uncertainty, (2) action uncertainty, (3) observation uncertainty, and (4) mutual information between the query and observation given the previous trajectory. For these estimations, we can leverage the output probability [malinin2021uncertainty, aichberger2024rethinking], consistency in multiple trials [farquhar2024detecting], world modeling [chae2025web, anonymous2026dreamphase], and a learning neural estimator [molavipour2020conditional, gritsai2025mist]. We leave the actual implementation for future work. See Appendix D for additional notes.
@@ 5 Practical Implications
Developing an agent UQ framework is not merely a theoretical exercise but a prerequisite for deploying LLM agents in non-deterministic real environments. We outline implications for frontier LLM research and three specialized domains: healthcare, coding, and robotics, in which agent UQ may have profound downstream effects, thereby incentivizing policymakers, practitioners, and researchers.
@@ 5.3 Reliable Software Engineering Agents
Although advanced coding agents [anthropic2025claude45systemcard, openai_gpt52codex2025] have pushed performance on standard benchmarks [jimenez2024swebench] to an impressive degree, the disparity between passing a controlled test and safely modifying a production-level codebase stresses the urgent need for an agent UQ framework. Agent UQ enables a more controllable exploration and commitment mechanism, wherein the agentic system knows uncertainty over bug patch candidates and decides whether to gather more evidence, e.g., check more files, or commit to a fix. Besides, considering uncertainty-triggered rollback and branching algorithm, in which an agent treats increasing uncertainty after an action as a signal to revert or branch by mirroring the checkpoint-rollback workflows of human engineers. These highlight the potential of UQ for reliable agentic coders.
5.4 Relible Embodied Agents in Cyber-Physical Systems
@@ 6 Open Problems
Although we have so far presented a conceptual roadmap and the practical implications of agent UQ, significant open problems remain as follows.
Developing rigid benchmarks at scale.
As in the case of process reward modeling mentioned in Section 3.2, annotating across long-horizon trajectories introduces non-trivial efforts and costs [choudhury2025process]. Moreover, as true uncertainty is unmeasurable, i.e., only a proxy-based evaluation is possible, establishing benchmarks for agent UQ stands on the next level of difficulty than the static LLM evaluation [ielanskyi2025addressing]. Investigation on defining proper agent performance measures [kapoor2025holistic, hendrycks2025definition] with scalable evaluation and annotation methods for agent rollouts [yang2024react, chang2024agentboard, kim2025beyond] would be a crucial stepping stone.
Long-horizon black-box uncertainty estimation.
LLM agents are mostly offered by proprietary services, where we cannot access the output probability of the agent. While model owners and providers can leverage low-cost white-box UQ approaches, others may have to rely on expensive black-box UQ that requires heavy sampling per action over a long trajectory [zhang2024luq]. This poses a great challenge for low-resource users in developing UQ with state-of-the-art agents.
Solution multiplicity.
Is high uncertainty per step due to the agent’s lack of knowledge about the right action for that step? Or, is it due to the intrinsic multiplicity of valid actions in that step? We cannot identify which of them is the right source of high uncertainty! Besides, in contrast to the classic question answering, the task bid to an agent is too complex to be perfectly specified at once. Thus, each intermediate step is intrinsically ambiguous and allows multiple valid actions, making it impossible to specify the source of the high uncertainty.
UQ in multi-agent or self-evolving agent systems.
While we confine our setup to single-agent, non-evolving agent systems, there has been remarkable advancement in multi-agent systems [guo2024large, wu2024autogen] and self-evolving agents [wang2024voyager, jiang2025adaptation]. Modeling the joint uncertainty dynamics of multiple agents or considering the non-stationary toolkit and agent parameters over time brings extra challenges that may require deliberation and other formal toolkits [tsallis2004nonextensive, j2016generalising, cheng2021general].
@@ 7 Conclusion
We argue that LLM UQ research should move towards a more realistic setup, i.e., agentic inference, featuring the interactive long-horizon task episode, which requires a paradigm shift from a point-wise estimate to the bi-directional dynamics model of uncertainty. We laid the foundation for this by providing a concrete definition and general formulation of agent UQ that abstracts a wide variety of existing UQ approaches. To tackle the limitations of existing framing, we presented a novel perspective, conditional uncertainty reduction process, with actionable guidelines and a theoretical analysis. As a concluding remark, we quote Schrödinger’s note on living in physics [schrodinger1974what]: “Every living means an increase in entropy of the part of the world, and it can only survive by continually drawing negative entropy from its environment.”
@@
Although this paper provided a novel high-level vision for agent UQ with a crystallized formulation, we built this vision on a relatively simple setup with some assumptions: reliable observation $O_{i}$ from the environment and user, and deterministic environment update $E_{i}\rightarrow E_{i+1}$. However, agentic systems in the wild may encounter unreliable (even adversarial; kang2025trap) observations and a stochastically evolving environment, where our current proposal does not ensure its effectiveness. In that scenario, we may need to explicitly model uncertainty in the agent’s evidence stream as well, e.g., trust/ignorance about tool outputs, retrieval, and user-provided facts, by leveraging formalisms, such as subjective logic [jsang2018subjective, cheng2020there], imprecise probability theory [walley1991statistical], and formal belief representation [formalbelief22], to safely manage memory and act under unreliable observations in stochastic world.
@@ Acknowledgments
We sincerely thank Artem Shelmanov, Shawn Im, Hyeong Kyu Choi, Jongwon Jeong, Eunsu Kim, Mingyu Kim, Ayoung Lee, JungEun Kim, Hayun Lee, and Kyungwoo Song for their sharp feedback on the draft. This work is supported in part by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation under awards IIS-2237037 and IIS-2331669, Office of Naval Research under grant number N00014-23- 1-2643, Schmidt Sciences Foundation, Open Philanthropy, Alfred P. Sloan Fellowship, and gifts from Google and Amazon. Paul Bogdan acknowledges the support by the National Science Foundation (NSF) under the NSF Award 2243104 under the Center for Complex Particle Systems (COMPASS), NSF Mid-Career Advancement Award BCS-2527046, U.S. Army Research Office (ARO) under Grant No. W911NF-23-1-0111, Defense Advanced Research Projects Agency (DARPA) Young Faculty Award and DARPA Director Fellowship Award, Okawa foundation award, National Institute of Health (NIH) under R01 AG 079957, and Intel faculty awards.
@@
We cover the following content in this Appendix.
  • (A) Additional Details on Formulation
  • (B) Alternative Uncertainty Measures
  • (C) Extended Context
  • (D) Note on Implementation
  • (E) Action Categorization
  • (F) Missing Proof
  • (G) Discussion and Future Work
Appendix A Additional Details on Formulation
A.1 Expression of various agentic prompting methods under our formulation.
In Section 3.1, we have introduced a definition of stochastic agent system (Def. 1) with a probabilistic graphical model of specific conditional dependency assumptions 2. Under this model, we can abstract some representative agentic prompting methods [ahn2022can, yao2022react, shinn2023reflexion, patil2025berkeley] with a unified language yet distinctive conditions as below.

### v2

@@ 3.2 A Unified View on Existing UQ Setups
We further draw a connection between this multi-turn UQ problem and the process reward modeling problem [uesato2022solving, lightman2024lets, li2025pqm] where we assign credit (reward) for each intermediate reasoning step, and then it can be utilized to evaluate the reasoning trajectory by aggregating the rewards across steps to derive a trajectory-level reward. For example, lightman2024lets proposed the product of the step-wise probabilities and the minimum probability between steps as the aggregation method. These are analogs to Eq. 4 and Eq. 5, respectively.
4 Technical Challenges in Agent UQ
We discuss why agentic setups are more challenging for UQ than non-agentic ones, identifying four distinct challenges. To motivate discussion, we present preliminary results on $\tau^{2}$-bench using existing UQ methods: negative log-likelihood (NLL), entropy over token probabilities (Entropy), and verbalized confidence, each aggregated over all tokens or turns generated by an agent throughout the trajectory (Appendix F has details).
4.1 Selection of Uncertainty Estimator
Approach Access Inference Overhead Theoretical Ground
Probability-based Free
Consistency-based Significant
Verbalized Confidence Neglectable
Table 1: Comparison between UQ approaches. Existing UQ approaches have unique weaknesses that become severe obstacles in agentic setups.
So what kind of uncertainty estimator should we adopt to model the agent’s uncertainty, $U(\cdot)$ in Eq. 3.1? We can start by examining three representative approaches that have been advanced in LLM UQ literature so far: output probability-based (and/or internal hidden state-based), response consistency-based, and prompting-based verbalized confidence methods. These approaches involve clear trade-offs along three dimensions: accessibility, inference-time cost overhead, and theoretical grounding. Probability-based methods exhibit limited applicability due to their reliance on access to output probabilities. While consistency-based and verbalized confidence methods do not face this accessibility constraint, they incur substantial inference-time overhead from repeated generation and lack theoretical grounding, respectively.
Scenario Avg. $r$ NLL Entropy Verb. Conf.
AUROC $\rho$ $\tau$ AUROC $\rho$ $\tau$ AUROC $\rho$ $\tau$
GPT-4.1 on Retail 0.509 0.597 0.169 0.138 0.580 0.139 0.114 0.575 0.179 0.170
GPT-4.1 on Telecom 0.517 0.624 0.214 0.176 0.611 0.192 0.158 0.685 0.330∗∗∗ 0.286∗∗∗
Kimi-K2.5 on Retail 0.447 0.469 -0.054 -0.044 0.468 -0.056 -0.045 0.523 0.039 0.032
Kimi-K2.5 on Telecom 0.965 0.645 0.093 0.076 0.664 0.104 0.086 0.580 0.051 0.042
Table 2: Evaluation of uncertainty estimators on $\tau^{2}$-bench. NLL and Entropy predict task failure (1r1-r); Verbalized confidence predicts task success ($r$), where $r$ denotes the reward; $\rho$ and $\tau$ denote Spearman and Kendall rank correlation coefficient, respectively. The significance is noted as $^{*}p<0.05$ and $^{***}p<0.001$.
Although this analysis applies to the classical single-turn QA setup as well, agentic scaffold turns these weaknesses into much more serious challenges: (1) probability-based methods cannot be applied to most frontier LLMs that centered at challenging agent benchmarks, as they do not consistently provide probability outputs; moreover, free-form long generation makes aggregated token probabilities less informative; (2) consistency-based methods become infeasible due to their prohibitively high inference cost in long-horizon, multi-turn settings; and (3) dynamically expanding context memory of agent (i.e., $E_{i}$ in Eq. 3.1), often coupled with noisy observations, results in increasingly inflated and unreliable verbalized confidence. A pilot study in Table 2 provides some results of probability-based and verbalized confidence methods for estimating the trajectory uncertainty $U(\mathcal{F}_{\leq T})$ with aggregated action uncertainty $\frac{1}{T}\sum_{i=1}^{T}U(A_{i})$. In many cases, the considered methods show close-to-random classifier performance in predicting the agent’s success/failure. We believe that establishing a theoretically grounded verbalized confidence method is a promising future direction, given its vivid merits in agentic setups such as accessibility and minimal extra inference cost.
4.2 Uncertainty of Heterogeneous Entities
Figure 3: Observation uncertainty on $\tau^{2}$-bench. We compare distributions over observation uncertainty estimates, i.e., average NLL over all user messages, from the agent LLM and the ground truth user simulator LLM.
Inference of agents involves not only the actions $A_{i}$ from the agent itself, but more importantly, observations $O_{i}$ through the conversations with the user and interactions with the environment via tools. There are no trivial means to estimate the uncertainty of observations that come from these heterogeneous external entities, departing from the epistemic uncertainty of the agent-self.
One may argue that we can just feed the whole trajectory, including the observation strings from the user and tools, into the agent and then use its output probability $P_{\pi,\mathcal{T}}(O_{i}|A_{i},E_{i})$ to approximate the underlying true observation distribution $P(O_{i}|A_{i},E_{i})$ per turn and corresponding uncertainty $U(O_{i}|A_{i},E_{i})$. However, the distribution of human language is definitely different from that of LLMs [munozortiz2024contrasting, alex2025do], and the distribution over the tool-calling outcome strings might be different too. To quantify this gap, we conduct a preliminary experiment on $\tau^{2}$-bench, where the user is simulated with another LLM equipped with a distinct prompt from that of the agent LLM. In Figure 3, we compare the trajectory-level average NLL density plots driven from the ground truth user simulator and the agent LLM. We see remarkable deviations between them, stressing the need for deliberation to estimate observation uncertainty. Adopting auxiliary LLM as an approximated world model can be a possible solution, where we see some promises in Figure 8.
4.3 Modeling Uncertainty Dynamics in Interactive Systems
One of the most distinctive features of agent UQ is that the model now operates in a non-isolated environment to generate a trajectory from iterative interactions. This presents a unique challenge: modeling multi-turn cascade uncertainty in an interactive system. Here, the desired UQ method should not only model the propagation of uncertainty across the trajectory but also account for additional information gained through interactions. To be specific, an agent in this open environment can gather extra information, make clarifications, and request confirmations at each step by interacting with the user and databases through conversations and tool-calling. These interactive actions are the driving force of goal achievement in the complex (and usually under-specified) agentic tasks in the wild [li2026thinking]. Therefore, the uncertainty model, which can reliably inform the agent’s failures in this setting, should reflect conditional uncertainty reduction through information-seeking behavior as well as the uncertainty propagation.
Figure 4: Uncertainty evolution across the trajectory on $\tau^{2}$-bench. We visualize the evolution of action uncertainty estimates across the normalized turn index, averaged over the entire trajectories from two groups: task success and failure.
Unfortunately, most existing multi-step UQ approaches discussed from Eq. 3 to Eq. 6 do not consider the interactivity of actions for each turn, and only focus on uncertainty propagation, while neglecting the reducible nature of uncertainty dynamics in non-isolated, interactive systems. In Figure 4, we present the uncertainty evolution of a naive averaging-based uncertainty aggregation method (Eq. 6), which is averaged over all trajectories from both success and failure groups. This simple, action-independent aggregation fails to meaningfully discern the failure group from the success group; Even the failure group achieves a sharper uncertainty decrease in the later part of trajectories in Telecom domain. This implies that the action-independent naive cascade modeling cannot robustly identify the agent’s failure both in the process and at the termination of the task, indicating the necessity of a new framework, pursuing explicit action-conditional modeling. We discuss a possible recommendation on this action-conditional uncertainty dynamics model in Appendix LABEL:sec:adpx:action-conditional.
4.4 Lack of Fine-grained Benchmarks
In the meantime, similar to the process reward modeling [choudhury2025process], annotating long-horizon agent trajectories incurs non-trivial effort and cost. The absence of fine-grained benchmarks becomes a natural bottleneck to developing a solid UQ method for agents. In this subsection, we provide a mini-survey of existing agent or multi-turn interaction benchmarks (See Appendix G), focusing primarily on LLM-centric agents rather than multimodal or physical ones. Specifically, we categorize 44 benchmarks into three classes depending on evaluation granularity: trajectory-level (conducted once at the end of the trajectory), milestone-level (involving several intermediate milestones or events), and turn-level (conducted at every single turn). Figure 5 reveals that only 4 out of 44 benchmarks provide turn-level annotation, which is very scarce compared to the trajectory-level benchmarks. Besides, although the milestone-level alternative shows better population than the turn-level one, it is still lacking, resulting in three times less than the trajectory-level one.
Figure 5: Survey on 44 modern agent benchmarks. The number denotes the percentage of papers in each category.
There are several obstacles to developing fine-grained benchmarks for agentic inference setups. First, agent benchmarks often involve very long rollouts that can easily exceed a hundred steps, making it hard to manually annotate each step. Although using frontier models as a substitute for human labeling has become increasingly common, running such models on long-horizon tasks can be prohibitively expensive for many researchers. Moreover, agent benchmarks are usually defined by complex policies; even for humans, it may take a couple of hours to fully understand the task configuration. To accelerate the development of agent UQ, we should pay greater attention to scalable dataset construction strategies for agentic inference.
@@ 5 Practical Implications
Developing the agent UQ framework is not merely a theoretical exercise but a prerequisite for deploying LLM agents in non-deterministic real environments. We outline implications for frontier LLM research and three specialized domains: healthcare, programming, and robotics, in which agent UQ may have profound downstream effects, thereby incentivizing policymakers, practitioners, and researchers.
@@ 5.3 Reliable Software Engineering Agents
Although advanced coding agents [anthropic2025claude45systemcard, openai_gpt52codex2025] have pushed performance on standard benchmarks [jimenez2024swebench] to an impressive degree, the disparity between passing a controlled test and safely modifying a production-level codebase stresses the urgent need for a reliable agent UQ framework. Agent UQ enables a more controllable exploration and commitment mechanism, wherein the agentic system knows uncertainty over bug patch candidates and decides whether to gather more evidence, e.g., check more files, commit to a fix, or interact with the user. Besides, considering uncertainty-triggered rollback and branching algorithm, in which an agent treats increasing uncertainty after an action as a signal to revert or branch by mirroring the checkpoint-rollback workflows of human engineers. These highlight the potential of UQ for reliable agentic coders.
5.4 Reliable Embodied Agents in Cyber-Physical Systems
@@ 6 Open Problems
We have so far presented a roadmap for agent UQ with a general formulation, identified challenges, and practical implications, but important open problems remain as follows.
Intrinsic solution multiplicity.
Is high uncertainty per step due to the agent’s lack of knowledge about the right action for that step? Or, is it due to the intrinsic multiplicity of valid actions in that step? We cannot identify which of them is the right source of high uncertainty! Besides, in contrast to the classic question answering, the task bid to an agent is too complex to be perfectly specified at once. Thus, each intermediate step is intrinsically ambiguous and allows multiple valid actions, making it impossible to specify the source of the high uncertainty. Graph-based parallelism/multiplicity modeling might be worth exploring as a possible solution here [wu2025regular, choi2026modex].
Evaluation beyond task failure.
Evaluation of predictive uncertainty has been conducted by measuring the informativeness of uncertainty estimates when predicting the incorrectness of the final answer. While suitable in single-turn QA settings, it becomes insufficient when we move to agentic LLM setups. In such setups, the sole prediction of eventual task failure collapses a rich, structured problem-solving process into a single scalar judgment, discarding useful information that makes uncertainty more actionable. We envision a new evaluation protocol that integrates dynamic solution multiplicity, task difficulty, and irreducible external ambiguity, as well as task failure, that may redefine uncertainty from a multi-facet view [guo2025uncertaintyprofiles]. This reframing calls for new metrics that credit not just predictive correlation with outcomes but the informativeness of uncertainty signals in multiple factors in agentic scaffolding, which is yet to be explored.
Uncertainty modeling in multi-agent systems.
Although we confine our attention to a single-agent environment, there has been notable advancement in multi-agent systems (MAS) [Liang2023EncouragingDT, du2024improving, guo2024large, wu2024autogen, choi2025debatevote]. Uncertainty dynamics in this MAS can be leveraged to improve the quality of inter-agent communication [yoffe2025debunc] and mitigate debate collapse [tang2026value], thus worth exploring. Unfortunately, while there are some works that borrow the strength of collective intelligence from MAS to improve UQ quality in a single-turn QA setup [yang2024confidence, feng2025rethinking], no existing work in this field attempts to design a UQ method tailored for MAS. Modeling the joint uncertainty dynamics of multiple agents brings extra challenges that may require deliberation and other formal toolkits [tsallis2004nonextensive, j2016generalising, cheng2021general].
Uncertainty modeling in self-improving agents.
In our setup, we assume that the agent only operates within a single episode. However, the agent research community has recently shown an increasing interest in self-evolving/improving agents [bai2022constitutional, shinn2023reflexion, NEURIPS2023_91edff07, wang2024voyager, jiang2025adaptation] that adapt across multiple episodes. Uncertainty information can also be utilized in this setup for reliable continual adaptation [zhang2026selaur]. However, the interactive self-evolving environment adds another non-trivial factor that shifts uncertainty dynamics as shown in huang2025beyond. Future exploration in this setup should reflect the non-stationary toolkit, context memory, and agent model parameters across multiple episodes.
@@ 7 Conclusion
We argue that LLM UQ research should move towards a more realistic setup, i.e., agentic inference, featuring the interactive long-horizon task episode, which requires a paradigm shift from a point-wise estimate to the sequential dynamics model of uncertainty. We laid the foundation for this by providing a concrete definition and general formulation of agent UQ that abstracts a wide variety of existing UQ approaches. With that, we identified four key challenges emerging in UQ on agentic scaffolding with detailed explanations and numerical analyses. We further note the future direction to explore agent UQ from the perspectives of real-world applications and open problems to provoke forward-looking discussions. We hope the three pillars provided in this work help communities design their upcoming works, products, and more.
@@
Agentic systems in the wild may encounter unreliable (even adversarial; kang2025trap) observations and a stochastically evolving environment. However, we do not dive deeper on that settings, and our formulation might not provide a proper abstraction. In that scenario, we may need to explicitly model uncertainty in the agent’s evidence stream as well, e.g., trust/ignorance about tool outputs, retrieval, and user-provided facts, by leveraging formalisms, such as subjective logic [jsang2018subjective, cheng2020there], imprecise probability theory [walley1991statistical], and formal belief representation [formalbelief22], to safely manage memory and act under unreliable observations in stochastic world.
@@ Acknowledgments
We sincerely thank Artem Shelmanov, Shawn Im, Hyeong Kyu Choi, Jongwon Jeong, Eunsu Kim, Mingyu Kim, Ayoung Lee, JungEun Kim, Hayun Lee, Hoyoon Byun, and Kyungwoo Song for their sharp feedback on the draft. This work is supported in part by the AFOSR Young Investigator Program under award number FA9550-23-1-0184, National Science Foundation under awards IIS-2237037 and IIS-2331669, Office of Naval Research under grant number N00014-23- 1-2643, Schmidt Sciences Foundation, Open Philanthropy, Alfred P. Sloan Fellowship, and gifts from Google and Amazon. Paul Bogdan acknowledges the support by the National Science Foundation (NSF) under the NSF Award 2243104 under the Center for Complex Particle Systems (COMPASS), NSF Mid-Career Advancement Award BCS-2527046, U.S. Army Research Office (ARO) under Grant No. W911NF-23-1-0111, Defense Advanced Research Projects Agency (DARPA) Young Faculty Award and DARPA Director Fellowship Award, Okawa foundation award, National Institute of Health (NIH) under R01 AG 079957, and Intel faculty awards.
@@
Appendix A Revision History
A prototype version of this paper titled "Towards Reducible Uncertainty Modeling for Reliable Large Language Model Agents" was released on arXiv in early February 2026. It was specifically targeted at the third challenge described in the current Section 4.3 by proposing a conceptual framework, conditional uncertainty reduction process, to enable action-conditional uncertainty dynamics modeling for interactive agents, which is available now at Section E in this Appendix. After receiving some early feedback from the community, we revised this article to cover a broad range of practical challenges emerging under agentic scaffolding, supported by quantitative analyses, rather than focusing narrowly on a single challenge and its corresponding solution. In the main body of the paper, Section 4 has undergone a major revision, while the remaining sections were subject to only moderate or minor edits.
Appendix B Additional Details on Formulation
B.1 Expression of Various Agentic Prompting under Our Formulation
In Section 3.1, we have introduced a definition of stochastic agent system (Def. 1) with a probabilistic graphical model of specific conditional dependency assumptions in Figure 2. Under this model, we can abstract some representative agentic prompting methods [ahn2022can, yao2022react, shinn2023reflexion, patil2025berkeley] with a unified language yet distinctive conditions as below.